# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v3

@@ Abstract
Modeling movement in real-world tasks is a fundamental goal for motor control, biomechanics, and rehabilitation engineering. However, widely used data-driven models of essential tasks like locomotion make simplifying assumptions such as linear and fixed timescale mappings between past inputs and future actions, which do not generalize to real-world contexts. Here, we develop a deep learning-based framework for action prediction with architecture-dependent trial embeddings, outperforming traditional models across contexts (walking and running, treadmill and overground, varying terrains) and input modalities (multiple body states, gaze). We find that neural network architectures with flexible input history-dependence like GRU and Transformer perform best overall. By quantifying the model’s predictions relative to an autoregressive baseline, we identify context- and modality-dependent timescales. These analyses reveal that there is greater reliance on fast-timescale predictions in complex terrain, gaze predicts future foot placement before body states, and the full-body state predictions precede those by center-of-mass-relevant states. This deep learning framework for action prediction provides quantifiable insights into the control of real-world locomotion and can be extended to other actions, contexts, and populations.
@@ Methods
Our framework is versatile and can be applied to any dataset with time-series inputs and discrete output events. To demonstrate this flexibility, we utilized several existing datasets of human locomotion across a range of contexts, including walking and running on both treadmills and overground, as well as across uneven terrain. The treadmill walking dataset was collected by Wang and Srinivasan [7], while the treadmill running dataset on even and uneven terrains was collected by Voloshina and colleagues [37]. The overground walking dataset, collected by Matthis and colleagues [30], includes full body kinematics of subjects walking on flat, medium, and rough terrains alongside gaze information recorded to track ground fixations during locomotion.
@@
We hypothesized that different input modalities are integrated throughout locomotion to inform future foot placements. These modalities include the CoM-relevant kinematics, the swing-foot kinematics, the full-body kinematics, and the gaze fixations (Figure 1A). The output predicted by the models consists of the lateral and fore-aft foot placement relative to the contact location of the opposite foot at heel strike (Figure 1B). In this study, we leveraged deep learning models to evaluate the relationships between these input modalities and the future foot placement. Let $v_{i}\in\{0,1,\dots,T-1\}$ be the trial ID of data point $i$, where T is the number of trials, and $l_{i}\in\left\{0,1\right\}$ be the left-right (L/R) flag of data point $i$, representing whether the corresponding output data point is a left (0) or a right (1) heel strike. The time series capturing the input modalities is denoted by $\mathbf{S}_{i}=\left[\mathbf{s}_{i}^{(1)},\mathbf{s}_{i}^{(2)},\dots,\mathbf{s}_{i}^{(m)}\right]$, where m is the number of features, and $\mathbf{s}_{i}^{(j)}$ is the $j$-th feature of the time series. This time series starts 6 steps prior to the predicted heel strike (Figure 1C) and its end is sampled from the previous gait cycle with 20 equally spaced gait phases. This range is chosen because 6 steps is a reasonable upper bound of how much subjects plan their foot placement ahead, while 2 steps is a reasonable lower bound [7, 2, 30]. Let $\phi$ denote the gait phase relative to one gait cycle prior to heel strike (i.e. previous heel strike of the same foot). With the dataset sampled at 20 time steps per gait cycle, the length of $\mathbf{s}_{i}^{(j)}$ ranges from 41 (when time series ends at $\phi=0$) to 61 (when time series ends at $\phi=1$). We define the concatenated inputs as $X=\{x_{i}\}_{i=1}^{n}$, where $x_{i}=\{\mathbf{S}_{i},v_{i},l_{i}\}$, and seek to obtain the mapping between those inputs $X$ and the outputs $Y=\{y_{i}\}_{i=1}^{n}$, where $y_{i}=\left[f_{i}^{ML},f_{i}^{AP}\right]$ where $f_{i}^{ML}$ and $f_{i}^{AP}$ represent the mediolateral (ML) and anteroposterior (AP) foot placement, respectively.
Figure 1: Overview of the data-driven modeling framework. A Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing-foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. B Foot placement output is defined by the lateral and fore-aft foot placement relative to the opposite foot at its previous heel strike. C The input is a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. D The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to provide additional trial-specific information. The incorporation of this embedding is architecture-dependent, i.e. it influences the model prediction differently depending on the neural network architecture, as detailed in the Methods section.
@@
Training deep learning models typically requires large amounts of data [40], which is challenging with locomotion as most datasets only capture a few minutes of data for each trial. Therefore, it is often impractical to train deep learning models tailored to specific subjects or trials. To address this limitation, we pooled the data from all trials and train a unified model that includes trial ID as an input. By doing so, the model can learn trial-specific variations in a data-driven way while efficiently leveraging the entire dataset (Figure 1D). We first reshaped the left-right flag $l_{i}$ to match the time series input’s dimension and concatenated them, which resulted in a matrix of shape $(t,m+1)$, where $m$ is the number of features in the time series. The trial ID was embedded [41] into a vector $E$ of length $\left|E\right|=\lceil\sqrt{T}\rceil$ and integrated with the time series data in a manner that is dependent on how each neural network processes the timesteps. The embedding enables us to isolate the inter-trial variability from the context- and modality-specific variability, which are the focuses of this study. The models we used include:
@@
Figure 2: Identification of key timepoints for dynamics and prediction. A The relative RMSE between modality-based predictions and the autoregressive baseline prediction as a function of gait phases. The timing of the peak difference indicates the phase where the input modality is the strongest predictor. B An example velocity curve along the locomotor axis, with the 5% of peak velocity threshold used to determine the onset of the swing phase. C The baseline $R^{2}$ curve, showing a breakpoint where $R^{2}$ increases significantly. The timing of this breakpoint is identified as the phase where the slope exceeds the average slope.
@@
Results
In this section, we present findings using our deep learning framework for action prediction to understand the influence of the environmental context and input modality on future foot placement. First, we demonstrate that neural network models, particularly GRUs, significantly outperform linear models in more challenging contexts. During overground walking, we find that subjects exhibit a tradeoff between fast and slow timescale predictions, with the magnitude of the tradeoff varying across different terrains. By comparing the predictive power of input modalities to an autoregressive baseline model, we reveal that control timescales are both modality- and context-dependent. Finally, we discover a gait phase mid-swing when the swing foot dynamics becomes useful for foot placement prediction.
Task Terrain M LSTM GRU FCNN TCN Trans LI LH LI2 LH2
TW N/A C 1.00 1.00 0.98 0.71 0.98 0.97 0.26 0.97 0.99
C 0.98 1.00 0.96 0.92 0.98 1.00 0.34 1.00 0.97
Even F 1.00 1.00 0.89 0.82 0.94 0.38 0.62 1.00 0.78
C 0.98 1.00 0.96 0.91 0.97 0.97 0.31 0.97 0.97
TR Uneven F 1.00 1.00 0.93 0.91 0.96 0.35 0.63 0.99 0.76
C 0.96 1.00 0.95 0.94 0.99 0.91 0.70 0.95 0.88
F 0.97 0.99 0.94 0.95 1.00 0.52 0.76 0.87 0.77
Flat V 0.80 0.87 0.70 0.91 1.00 0.69 0.30 0.75 0.30
C 0.96 0.99 0.95 0.98 1.00 0.86 0.66 0.93 0.78
F 0.97 0.98 0.94 0.99 1.00 0.52 0.76 0.82 0.77
Medium V 0.95 0.98 0.88 1.00 0.97 0.68 0.32 0.75 0.33
C 0.99 1.00 0.97 0.99 1.00 0.31 0.68 0.87 0.73
F 0.99 0.99 0.94 0.98 1.00 0.58 0.74 0.74 0.74
OW Rough V 1.00 1.00 0.92 0.96 0.97 0.34 0.61 0.61 0.61
Fast- versus slow-timescale predictions are context-dependent
Figure 3: Relationship between modality-based intercept and peak relative predictive power. A Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. B Baseline intercept versus CoM-relevant kinematics intercepts for treadmill walking. C Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. D Coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between long- and short-timescale prediction.
Figure 4: Predictive power across input modalities and contexts. A Model predictions for lateral foot placement using CoM-relevant kinematics during treadmill walking, shown across increasing input time windows. B Explained variance ($R^{2}$, median and interquartile range) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) prediction. The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. C Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. D Relative predictive power of CoM-relevant kinematics for treadmill walking. E Relative predictive power of CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
The swing foot dynamics predict future foot placement mid-swing
Figure 5: Timescale of swing foot prediction dynamics across contexts. A Timing of when the swing foot starts to swing across different contexts. B, C Comparison of the timing when the swing foot starts to swing with the timing of when the swing foot begins to provide significant information about foot placement, detected using the average slope method (Figure 2C), during treadmill (B) and overground (C) movements. ∗∗ for $p<0.01$; $n.s.$ for not significant.
Discussion
Tradeoffs between long- and short-timescale action predictions.
Input modality-dependent timescales for motor control.
Conclusions and future work.

### v4

@@ Abstract
Modeling human movement in real-world tasks is a fundamental goal for motor control, biomechanics, and rehabilitation engineering. However, existing models of essential tasks like locomotion are not applicable across varying terrain, mechanical conditions, and sensory contexts. This is at least in part due to simplifying assumptions like linear and fixed timescales mappings between inputs and future actions, which may not be broadly applicable. Here, we develop a deep learning-based framework for action prediction, outperforming traditional models across multiple contexts (walking and running, treadmill and overground, varying terrains) and input modalities (multiple body states, visual gaze). We find that neural network architectures with flexible input history-dependence, like GRU and Transformer, and with architecture-dependent trial embeddings perform best overall. By quantifying the model’s predictions relative to an autoregressive baseline, we identify context- and modality-dependent timescales. These analyses reveal that there is greater reliance on fast-timescale predictions in complex terrain, gaze predicts future foot placement before body states, and the full-body state predictions precede those by center-of-mass states. This deep learning framework for human action prediction provides quantifiable insights into the control of real-world locomotion and can be extended to other actions, contexts, and populations.
@@ Introduction
Results
In this section, we present findings using our deep learning framework for action prediction to understand the influence of the environmental context and input modality on future foot placement. First, we demonstrate that neural network models, particularly GRUs, significantly outperform linear models in more challenging contexts. During overground walking, we find that subjects exhibit a tradeoff between fast and slow timescale predictions, with the magnitude of the tradeoff varying across different terrains. By comparing the predictive power of input modalities to an autoregressive baseline model, we reveal that control timescales are both modality- and context-dependent. Finally, we discover a gait phase mid-swing when the swing foot dynamics becomes useful for foot placement prediction. We encourage readers to first read through the Methods section to understand the terminology used below. An overview of the key methodological components in the deep learning framework can be found in Figure 1.
Figure 1: Overview of the deep learning framework. A We include diverse tasks and contexts, including overground walking with various terrain conditions, treadmill walking, and treadmill running on even and uneven terrain. B Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing-foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. C Foot placement output is defined by the lateral and fore-aft foot placement relative to the opposite foot at its previous heel strike. D The input is a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. E The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to provide additional trial-specific information. The incorporation of this embedding is architecture-dependent, i.e. it influences the model prediction differently depending on the neural network architecture, as detailed in the Methods section.
Task Terrain M LSTM GRU FCNN TCN Trans LI LH LI2 LH2
TW N/A C 1.00 1.00 0.98 0.71 0.98 0.97 0.26 0.97 0.99
C 0.98 1.00 0.96 0.92 0.98 1.00 0.34 1.00 0.97
Even F 1.00 1.00 0.89 0.82 0.94 0.38 0.62 1.00 0.78
C 0.98 1.00 0.96 0.91 0.97 0.97 0.31 0.97 0.97
TR Uneven F 1.00 1.00 0.93 0.91 0.96 0.35 0.63 0.99 0.76
C 0.96 1.00 0.95 0.94 0.99 0.91 0.70 0.95 0.88
F 0.97 0.99 0.94 0.95 1.00 0.52 0.76 0.87 0.77
Flat V 0.80 0.87 0.70 0.91 1.00 0.69 0.30 0.75 0.30
C 0.96 0.99 0.95 0.98 1.00 0.86 0.66 0.93 0.78
F 0.97 0.98 0.94 0.99 1.00 0.52 0.76 0.82 0.77
Medium V 0.95 0.98 0.88 1.00 0.97 0.68 0.32 0.75 0.33
C 0.99 1.00 0.97 0.99 1.00 0.31 0.68 0.87 0.73
F 0.99 0.99 0.94 0.98 1.00 0.58 0.74 0.74 0.74
OW Rough V 1.00 1.00 0.92 0.96 0.97 0.34 0.61 0.61 0.61
Fast- versus slow-timescale predictions are context-dependent
Figure 2: Relative predictive power reveals fast and slow timescales across contexts. A Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. B Baseline intercept versus CoM-relevant kinematics intercepts for treadmill walking. C Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. D Coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between long- and short-timescale prediction.
Figure 3: Relative predictive power reveals modality-dependent control timescales. A Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. B Relative predictive power of CoM-relevant kinematics for treadmill walking. C Relative predictive power of CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
The swing foot dynamics predict future foot placement mid-swing
Figure 4: Swing foot dynamics predict future foot placement mid-swing. A Timing of when the swing foot starts to swing across different contexts. B, C Comparison of the timing when the swing foot starts to swing with the timing of when the swing foot begins to provide significant information about foot placement, detected using the average slope method (Figure 5C), during treadmill (B) and overground (C) movements. ∗∗ for $p<0.01$; $n.s.$ for not significant.
Discussion
Tradeoffs between long- and short-timescale action predictions.
Input modality-dependent timescales for motor control.
Conclusions and future work.
@@ Methods
Our framework is versatile and can be applied to any dataset with time-series inputs and discrete output events. To demonstrate this flexibility, we utilized several existing datasets of human locomotion across a range of contexts (Figure 1A), including walking and running on both treadmills and overground, as well as across uneven terrain. The treadmill walking dataset was collected by Wang and Srinivasan [7], while the treadmill running dataset on even and uneven terrains was collected by Voloshina and colleagues [37]. The overground walking dataset, collected by Matthis and colleagues [30], includes full body kinematics of subjects walking on flat, medium, and rough terrains alongside gaze information recorded to track ground fixations during locomotion.
@@
We hypothesized that different input modalities are integrated throughout locomotion to inform future foot placements. These modalities include the CoM-relevant kinematics, the swing-foot kinematics, the full-body kinematics, and the gaze fixations (Figure 1B). The output predicted by the models consists of the lateral and fore-aft foot placement relative to the contact location of the opposite foot at heel strike (Figure 1C). In this study, we leveraged deep learning models to evaluate the relationships between these input modalities and the future foot placement. Let $v_{i}\in\{0,1,\dots,T-1\}$ be the trial ID of data point $i$, where T is the number of trials, and $l_{i}\in\left\{0,1\right\}$ be the left-right (L/R) flag of data point $i$, representing whether the corresponding output data point is a left (0) or a right (1) heel strike. The time series capturing the input modalities is denoted by $\mathbf{S}_{i}=\left[\mathbf{s}_{i}^{(1)},\mathbf{s}_{i}^{(2)},\dots,\mathbf{s}_{i}^{(m)}\right]$, where m is the number of features, and $\mathbf{s}_{i}^{(j)}$ is the $j$-th feature of the time series. This time series starts 6 steps prior to the predicted heel strike (Figure 1D) and its end is sampled from the previous gait cycle with 20 equally spaced gait phases. This range is chosen because 6 steps is a reasonable upper bound of how much subjects plan their foot placement ahead, while 2 steps is a reasonable lower bound [7, 2, 30]. Let $\phi$ denote the gait phase relative to one gait cycle prior to heel strike (i.e. previous heel strike of the same foot). With the dataset sampled at 20 time steps per gait cycle, the length of $\mathbf{s}_{i}^{(j)}$ ranges from 41 (when time series ends at $\phi=0$) to 61 (when time series ends at $\phi=1$). We define the concatenated inputs as $X=\{x_{i}\}_{i=1}^{n}$, where $x_{i}=\{\mathbf{S}_{i},v_{i},l_{i}\}$, and seek to obtain the mapping between those inputs $X$ and the outputs $Y=\{y_{i}\}_{i=1}^{n}$, where $y_{i}=\left[f_{i}^{ML},f_{i}^{AP}\right]$ where $f_{i}^{ML}$ and $f_{i}^{AP}$ represent the mediolateral (ML) and anteroposterior (AP) foot placement, respectively.
@@
Training deep learning models typically requires large amounts of data [74], which is challenging with locomotion as most datasets only capture a few minutes of data for each trial. Therefore, it is often impractical to train deep learning models tailored to specific subjects or trials. To address this limitation, we pooled the data from all trials and train a unified model that includes trial ID as an input. By doing so, the model can learn trial-specific variations in a data-driven way while efficiently leveraging the entire dataset (Figure 1E). We first reshaped the left-right flag $l_{i}$ to match the time series input’s dimension and concatenated them, which resulted in a matrix of shape $(t,m+1)$, where $m$ is the number of features in the time series. The trial ID was embedded [75] into a vector $E$ of length $\left|E\right|=\lceil\sqrt{T}\rceil$ and integrated with the time series data in a manner that is dependent on how each neural network processes the timesteps. The embedding enables us to isolate the inter-trial variability from the context- and modality-specific variability, which are the focuses of this study. The models we used include:
@@
Figure 5: Identification of relevant timepoints. A The relative RMSE between modality-based predictions and the autoregressive baseline prediction as a function of gait phases. The timing of the peak difference indicates the phase where the input modality is the strongest predictor. B An example velocity curve along the locomotor axis, with the 5% of peak velocity threshold used to determine the onset of the swing phase. C The baseline $R^{2}$ curve, showing a breakpoint where $R^{2}$ increases significantly. The timing of this breakpoint is identified as the phase where the slope exceeds the average slope.